# Antidote: Post-fine-tuning Safety Alignment for Large Language Models against Harmful Fine-tuning

### v2

@@ Introduction
0:  Mask ratio, $\alpha$;Alignment dataset, $D_{align}$; Fine-tuning dataset, $D_{ft}$; Re-alignment dataset, $\mathcal{D}_{realign}$;
@@
  — — STAGE I (alignment) begins here — —
  $\check{\bm{w}}$ = SafetyAlignment($\bm{w}$) on $D_{align}$
  — — STAGE II (fine-tuning) begins here — —
  $\hat{\bm{w}}$ = Hamrfulfine-tuning($\check{\bm{w}}$) on $D_{ft}$
  — STAGE III (one-shot pruning) begins here
  Calculate importance score $h(\hat{\bm{w}},\mathcal{D}_{\text{realign}})$ with Eq. (1)
  $\hat{\bm{m}}$ = ArgTopKα($h(\hat{\bm{w}},\mathcal{D}_{realign})$)
  $\tilde{\bm{w}}=(\bm{1}-\tilde{\bm{m}})\odot\hat{\bm{w}}$
To achieve the defense goal, we first need to identify the important parameters over a particular dataset using Wanda score. The Wanda score (Sun et al. 2023) measures the importance score of parameters given a dataset $\mathcal{D}$, as follows:
$[h(\bm{w},\mathcal{D})]_{j}=\frac{1}{|\mathcal{D}|}\sum_{\bm{X}\in\mathcal{D}}|\bm{w}_{j}|\cdot\|\bm{X}_{j}\|_{2}$ (1)
where $[\cdot]_{j}$ retrieve the j-th element of the vector, $\bm{X}_{j}$ is the input (or activation) associated with the $j$-th weight coordinate, and $\bm{X}$ represents a data point in a given dataset $\mathcal{D}$.
To recognize the harmful parameters, one intuitive idea is to extract the topk most important mask (harmful mask) on the re-alignment dataset, indicating the most important parameters for harmful content generation, as follows.
$\hat{\bm{m}}=\text{ArgTopK}_{\alpha}(h(\hat{\bm{w}},\mathcal{D}_{realign}))$ (2)
Given the harmful mask $\tilde{\bm{m}}$ and the weights after fine-tuning $\hat{\bm{w}}$, The pruning operation is as follows:
$\tilde{\bm{w}}=(\bm{1}-\tilde{\bm{m}})\odot\hat{\bm{w}}$ (3)
where $\tilde{\bm{w}}$ is the re-aligned model that is ready for deployment.
In summary, at stage I, we align the model on the alignment dataset using a vanilla SFT solution. At stage II, we finetune the model on the user fine-tuning dataset with SFT, and obtain the harmful mask, which captures the important parameters for harmful data. Finally, at the post-fine-tuning stage, we apply the pruning mask to the model and recover it from the harmful behavior. See Algorithm 1 for full procedure.
@@
Experiments
Setup
Model and Datasets. We use three mainstream pre-trained models, i.e., Llama2-7B, Mistral-7B and Gemma-7B for evaluations. In the default setting, we use Llama2-7B as the backbone. We consider three datasets associated with harmful data. The first dataset is an alignment dataset, which contains alignment data (i.e., data paired with harmful prompt-safe answers). The second is fine-tuning the dataset. This dataset is mixed with $p$ (percentage) of harmful data (paired with harmful prompt-harmful answer) and $1-p$(percentage) of downstream data (e.g., SST2, GSM8K, etc). The last one is a re-alignment dataset, which is solely constituted by harmful data. The alignment data are sampled from BeaverTails (Ji et al. 2023) with the label is_safe=True. The harmful data in fine-tuning dataset and realignment are also sampled from BeaverTails (Ji et al. 2023) with is_safe=False, but the harmful data in those two datasets are different. For fine-tuning tasks, we consider four different datasets, i.e., SST2, AGNEWS, GSM8K and AlpacaEval. We discuss how to integrate and evaluate these tasks in supplementary materials.
Metrics. Following (Huang et al. 2024), we use two metrics for evaluation.
@@
  • Harmful Score (HS). We use the moderation model from (Ji et al. 2023) to flag the model output given unseen malicious instructions. Harmful score is the ratio of the flagged unsafe output.
To calculate the harmful score, we sample 1000 harmful instructions from BeaverTails (Ji et al. 2023). To calculate finetune accuracy, we sample 872, 1000, 1000, and 122 samples from the corresponding fine-tuning testing dataset.
Baselines. We mainly consider five baselines in evaluation. SFT is utilized to supervised fine-tuning on both the alignment and fine-tuning stages. Two representative alignment-stage solutions, i.e., Vaccine (Huang, Hu, and Liu 2024) and RepNoise (Rosati et al. 2024a) modify the alignment stage while keeping the fine-tuning stage optimization as SFT. Two representative fine-tuning-stage solutions, i.e., Lisa(Huang et al. 2024) and LDIFS (Mukhoti et al. 2023) modify the fine-tuning stage while keeping the alignment stage optimization as SFT.
Training Details and Hyper-parameters. We follow (Huang, Hu, and Liu 2024) to utilize LoRA (Hu et al. 2021) for alignment and fine-tuning. The rank of the adaptor is 256 for both tasks. For alignment, we use 5000 safety samples and an AdamW optimizer with a learning rate of 1e-3. We train the alignment data for 20 epochs. For fine-tuning, we use $n$ samples, among which $p$ (percentage) of samples are harmful data, an AdamW optimizer with a learning rate of $lr$ is used, and we train for $ep$ epochs. The default setting is $n=5000$, $p=0.2$, $lr=1e-4$ and $ep=20$, and the default dataset is SST2 unless otherwise specified. For hyper-parameters for Antidote, in default, we set the mask ratio to be $\alpha=0.2$ (specially, $\alpha=0.05$ for GSM8K) and we sample 2000 harmful samples to form the re-alignment dataset used by Antidote. See for a setting for hyper-parameters for baselines. All experiments are done with an H100.
@@
Table 4: Evaluation under different training epochs. The dataset is GSM8K and other settings are default.
@@
Robustness to fine-tuning samples. We show in Table 2 how different fine-tuning samples used in the fine-tuning stage will affect the practical defense performance. Our results indicate Antidote again achieves the best defense performance among the baselines with a remarkable 13.42% reduction of the harmful score. Antidote is again the only defense that is universally robust to different sample number settings.
Robustness to learning rate in fine-tuning. In Table 3, we adjust the learning rate in the fine-tuning stage to see its impact on defense performance. We use GSM8K for this evaluation to provide more diversified statistical data. Our results show that Antidote reduces 6.56% average harmful score with a marginal 0.38% average finetune accuracy drop. Although the harmful score reduction is less competitive compared to two fine-tuning-stage defenses Lisa and LDIFS (which respectively achieve 7.72% and 9.50% harmful score reduction), we note that their performance gain comes with a drastic reduction of finetune accuracy. Moreover, the result when $lr=1e-3$ also coincides with the finding in our motivation section, that both Lisa and LDIFS suffer from hyper-parameter sensitivity. In contrast, Antidote is less susceptible to the exact setting of learning rate.
@@
Generalizations to datasets. In Table 5, we show the evaluation results for different datasets. Our results confirm that Antidote can be generalized to different fine-tuning tasks. On average, we show that Antidote reduce the harmful score by 11.75% with 3.08% of finetune accuracy loss. Here we did not specifically tune the mask ratio $\alpha$ for each dataset, and we will discuss later that the tradeoff between harmful score and finetune accuracy can be adjusted by this key hyper-parameter.
Table 5: Evaluation on different datasets.
@@
Generalizations to models. We show in Table 6 how different methods perform on different LLMs. Our results indicate that Antidote can be generalized to different model architectures. For Llama2-7B, Mistral-7B, and Gemma-7B, Antidote respectively achieve 11.6%, 20.0%, 22.5% reduction of harmful score with a minor reduction of 1.49%, 0.92%, and 1.72% finetune accuracy. Particularly, in terms of finetune accuracy, our results coincide with the bench-marking results that the rank of language ability of three models is Gemma-7B$>$Mistral-7B$>$Llama2-7B. In terms of reducing harmful score, our results seem to indicate that Antidote is more effective when the backbone model is stronger.
@@
Statistical and System Evaluation
@@
Harmful embedding drift. To justify the reason why Antidote is able to recover the model even with a large learning rate and training epoch, we follow (Huang, Hu, and Liu 2024) to measure the harmful embedding drift (HED), which tracks the L2 norm of the difference between the hidden embedding of the aligned model and that of the finetuned model over the same alignment data. We show in Figure 5 how different learning rates and training epochs affect this statistic. As shown, Antidote maintains the HED on a small scale, while the HED of other mitigation strategies escalates with the growth of learning rate and training epochs. Particularly, note that Antidote and SFT share the same two identical processes (and therefore the same HED) in their first two stages, but after the one-shot pruning in the post-fine-tuning stage, the HED of Antidote is significantly lower. This justifies that pruning out the identified harmful parameters can indeed recover the alignment knowledge preserved in the model.
In summary, our finding justifies that Antidote can minimize the hidden embedding drift over the alignment data by pruning some specific harmful coordinates, and thereby mitigating the harmful embedding drift issue mentioned in (Huang, Hu, and Liu 2024).
@@
Hyper-parameters Analysis and Ablation Study
@@
Table 9: Evaluation of Antidote with different ways to calculate Wanda score under different poison ratio $p$. Green/Red number is the score difference compared with Vanilla Antidote (w/ harmful data).
@@
Impact of the size of realignment dataset. Per Eq. (1), we calculate the harmful Wanda score with the average statistics over the realignment dataset (harmful dataset). It is interesting to see how different numbers of harmful samples included in the re-alignment dataset can affect defense performance. Our results in Table 10 show that the general trend is that a larger number of harmful samples are preferable in terms of reducing harmful score. This is understandable because a larger number of harmful samples can better reflect the real harmful distribution therefore resulting in a more precise identification of harmful parameters. Another finding is that 1k samples of harmful data seem to be performing well, and the benefit of further increasing the sample number diminishes. As collecting 1k harmful samples is not too restrictive [^2] We have much larger scale harmful datasets available today. For example, BeaverTail (Ji et al. 2023) contains 330k harmful samples., this experiment validates the feasibility of Antidote.
@@
Visualization
As follows, we show how different methods respond to a malicious prompt. We used the finetuned model over the default setting for evaluation. As shown, Antidote can give refusal answers to the sensitive questions while others cannot.
@@ Appendix A Experiment Setup
Detailed Setup
Training hyper-parameters. For the alignment stage, we use the learning rate of $31e-$ 3 and train for 20 epochs. For the fine-tuning stage, we use the default learning rate $41e-$ 4 and train for 20 epochs in default. We follow (Huang, Hu, and Liu 2024) to utilize the double LoRA implementation, i.e., for the alignment stage and fine-tuning stage, two different LoRA adaptors are used. The rank of LoRA adaptor is 256 with LoRA alpha set to 4. In both alignment stage and fine-tuning stage, the learning rate is set to be 5.
Prompt template. We follow (Huang, Hu, and Liu 2024) to use the Alpaca prompt template (Taori et al. 2023) in the following for constructing supervised dataset for alignment/finetuning.

### v3

@@ 4 Methodology
0: Mask ratio, $\alpha$; Re-alignment dataset, $\mathcal{D}_{realign}$; Safety alignment-broken fine-tuned model, $\bm{w}$ ;
@@
 Calculate importance score $h(\bm{w},\mathcal{D}_{\text{realign}})$ with Eq. (1)
[…]
$\tilde{\bm{w}}=(\bm{1}-\bm{m})\odot\bm{w}$
Identify harmful parameters. To achieve the defense goal, we first need to identify the important parameters (i.e., harmful parameter) over the re-alignment dataset using Wanda score. The Wanda score (Sun et al., 2023) measures the importance score of parameters given the re-alignment dataset $\mathcal{D}_{realign}$, as follows:
$h(\bm{w},\mathcal{D}_{realign})]_{j}=\frac{1}{|\mathcal{D}|}\sum_{\bm{x}\in\mathcal{D}_{realign}}|\bm{w}_{j}|\cdot\|\bm{A}_{j}(\bm{x},\bm{w})\|_{2}$ (1)
where $\mathcal{D}_{realign}$ is the harmful dataset containing harmful question-harmful answer pairs and $\bm{w}$ is a vector representing model weights of the safety-alignment broken fine-tuned model. $\cdot]_{j}$ retrieves the $j$-th element of the vector, $\bm{w}_{j}$ is the $j$-th weight coordinate (i.e., an element of the vector), $\bm{x}$ represents a data point in the re-alignment dataset $\mathcal{D}$, and $\bm{A}_{j}(\bm{x},\bm{w})$ retrieves the data point $\bm{x}$’s hidden activation that is associated with the $j$-th weight coordinate. Intuitively, the importance of a coordinate of parameter is related to its absolute value and the value of its input.
@@
To recognize the harmful parameters over the fine-tuned model weights $\bm{w}$, one intuitive idea is to extract the topk most important mask (harmful mask) on the re-alignment dataset, indicating the most important parameters for harmful content generation, as follows.
$\bm{m}=\text{ArgTopK}_{\alpha}(h(\bm{w},\mathcal{D}_{realign}))$ (2)
where $\text{ArgTopK}_{\alpha}()$ returns a mask with the topk coordinate being 1 and the rest being 0. $\alpha$ is the ratio of coordinates that are masked to 1, which we name mask ratio for simplicity.
Removal of Harmful parameter. Given the harmful mask $\bm{m}$ and the weights after fine-tuning $\bm{w}$, The pruning operation of the harmful parameters is as follows:
$\tilde{\bm{w}}=(\bm{1}-\bm{m})\odot\bm{w}$ (3)
where $\tilde{\bm{w}}$ is the re-aligned model that is ready for deployment, and $\odot$ is the Hadamard product, which multiples the two vectors for each element.
In summary, given a safety alignment-broken fine-tuned model, we first identify the top-k harmful parameters with a harmful mask. Then we remove those harmful parameters from the fine-tuned model to recover it from the harmful behavior. The recovered model is then deployed to serve users’ customized tasks. See Algorithm 1 for full procedure.
5 Experiments
5.1 Setup
Model and Datasets. We use three mainstream pre-trained models, i.e., Llama2-7B, Mistral-7B and Gemma-7B for evaluations. In the default setting, we use Llama2-7B as the backbone. We consider three datasets associated with harmful data. The first dataset is an alignment dataset, which contains alignment data (i.e., data paired with harmful prompt-safe answers). The second is fine-tuning the dataset. This dataset is mixed with $p$ (percentage) of harmful data (paired with harmful prompt-harmful answer) and 1p1-p(percentage) of downstream data (e.g., SST2, GSM8K, etc). The last one is a re-alignment dataset, which is solely constituted by harmful data. The alignment data are sampled from BeaverTails (Ji et al., 2023) with the label is_safe=True. The harmful data in fine-tuning dataset and realignment are also sampled from BeaverTails (Ji et al., 2023) with is_safe=False, but the harmful data in those two datasets are different. For fine-tuning tasks, we consider four different datasets, i.e., SST2, AGNEWS, GSM8K and AlpacaEval. We discuss how to integrate and evaluate these tasks in supplementary materials.
Metrics. Following (Rosati et al., 2024a; Hsu et al., 2024; Huang et al., 2024d, b, a), we use two metrics for evaluation. Both the two metrics are measured over the fine-tuned model (except for Antidote, they are measured over the re-alignment model).
@@ 5.1 Setup
  • Harmful Score (HS). We use the moderation model from (Ji et al., 2023) to flag the model output given unseen malicious instructions. Harmful score is the ratio of the flagged unsafe output.
To calculate the harmful score, we sample 1000 harmful instructions from BeaverTails (Ji et al., 2023). To calculate finetune accuracy, we sample 872, 1000, 1000, and 122 samples from the corresponding fine-tuning testing dataset. In testing time, we use greedy decoding for text generation.
Baselines. We mainly consider five baselines in evaluation. SFT is utilized to supervised fine-tuning on both the alignment and fine-tuning stages. Two representative alignment-stage solutions, i.e., Vaccine (Huang et al., 2024d) and RepNoise (Rosati et al., 2024a) modify the alignment stage while keeping the fine-tuning stage optimization as SFT. Two representative fine-tuning-stage solutions, i.e., Lisa(Huang et al., 2024b) and LDIFS (Mukhoti et al., 2023) modify the fine-tuning stage while keeping the alignment stage optimization as SFT.
Training Details and Hyper-parameters. We follow (Huang et al., 2024d) to utilize LoRA (Hu et al., 2021) for alignment and fine-tuning. The rank of the adaptor is 256 for both tasks. For alignment, we use 5000 safety samples and an AdamW optimizer with a learning rate of 1e-3. We train the alignment data for 20 epochs. For fine-tuning, we use $n$ samples, among which $p$ (percentage) of samples are harmful data, an AdamW optimizer with a learning rate of $lr$ is used, and we train for $ep$ epochs. The default setting is $n=5000$, $p=0.2$, lr=1e4lr=1e-4 and $ep=20$, and the default dataset is SST2 unless otherwise specified. For hyper-parameters for Antidote, in default, we set the mask ratio to be $\alpha=0.2$ (specially, $\alpha=0.05$ for GSM8K) and we sample 2000 harmful samples to form the re-alignment dataset used by Antidote. See for a setting for hyper-parameters for baselines. All experiments are done with an H100.
@@
Table 4: Evaluation under different fine-tuning epochs. The dataset is GSM8K and other settings are default.
@@
5.2 Main Results
Robustness to fine-tuning samples. We show in Table 2 how different fine-tuning samples used in the fine-tuning stage will affect the practical defense performance. Our results indicate Antidote again achieves the best defense performance among the baselines with a remarkable 13.42% reduction of the harmful score. Antidote is again the only defense that is universally robust to different sample number.
Robustness to benign fine-tuning attack. (Qi et al., 2023) and a few subsequent research (He et al., 2024; Guan et al., 2025) show that fine-tuning on benign data can also degrade the model’s safety alignment. Next we show in Table 5 that Antidote can also be robust to benign fine-tuning attack. As shown, Antidote can sufficiently decrease the harmful score but without hurting much fine-tune accuracy.
Table 5: Evaluation of benign fine-tuning attack on GSM8K.
Harmful Score Fine-tune Accuracy
SFT 61.50 27.60
RepNoise 66.10 27.40
Vaccine 58.90 26.60
LDIFS 64.40 6.70
Lisa 59.20 27.60
Antidote 57.10 27.80
Robustness to learning rate in fine-tuning. In Table 3, we adjust the learning rate in the fine-tuning stage to see its impact on defense performance. We use GSM8K for this evaluation to provide more diversified statistical data. Our results show that Antidote reduces 6.56% average harmful score with a marginal 0.38% average finetune accuracy drop. Although the harmful score reduction is less competitive compared to two fine-tuning-stage defenses Lisa and LDIFS (which respectively achieve 7.72% and 9.50% harmful score reduction), we note that their performance gain comes with a drastic reduction of finetune accuracy. Moreover, the result when lr=1e3lr=1e-3 also coincides with the finding in our motivation section, that both Lisa and LDIFS suffer from hyper-parameter sensitivity. In contrast, Antidote is less susceptible to the exact setting of learning rate.
@@ 5.2 Main Results
5.3 Generalizations on Datasets and Models
Generalizations to fine-tuning datasets. In Table 6, we show the evaluation results for different datasets. Our results confirm that Antidote can be generalized to different fine-tuning tasks. On average, we show that Antidote reduce the harmful score by 11.75% with 3.08% of finetune accuracy loss. Here we did not specifically tune the mask ratio $\alpha$ for each dataset, and we will discuss later that the tradeoff between harmful score and finetune accuracy can be adjusted by this key hyper-parameter.
Table 6: Evaluation on different fine-tuning datasets.
@@ 5.3 Generalizations on Datasets and Models
Generalization to alignment datasets. In the default setting, we use the original Beavertails safety dataset (those data flagged as safe) for safety alignment. Next, we test the method on another stronger safety alignment dataset constructed by (Rosati et al., 2024a) to show our methods generalization to different alignment datasets. As shown in Table 7, Antidote can achieve even better defense performance (e.g., over 40% of HS reduction when $p=0.2$) under safety alignment with stronger alignment dataset. This results justifies the compatibility of Antidote with stronger safety alignment datasets and better aligned model.
Table 7: Using BeaverTails refusal (Rosati et al., 2024a) as safety alignment dataset.
p=0 p=0.05 p=0.1 p=0.2 p=0.5
Methods HS FA HS FA HS FA HS FA HS FA
SFT 13.5 29.2 80.3 28.2 78.8 28.1 78.6 26.8 82.3 24.1
Lisa 45.5 29.7 67.8 28.5 75.5 28.5 78.7 27.2 78.7 24.1
Antidote 2.3 22.2 11.3 22.6 15.5 21.9 21.8 20.6 36.5 19.3
Generalizations to models. We show in Table 8 how different methods perform on different LLMs. Our results indicate that Antidote can be generalized to different model architectures. For Llama2-7B, Mistral-7B, and Gemma-7B, Antidote respectively achieve 11.6%, 20.0%, 22.5% reduction of harmful score with a minor reduction of 1.49%, 0.92%, and 1.72% finetune accuracy. Particularly, in terms of finetune accuracy, our results coincide with the bench-marking results that the rank of language ability of three models is Gemma-7B>>Mistral-7B>>Llama2-7B. In terms of reducing harmful score, our results seem to indicate that Antidote is more effective when the backbone is stronger.
@@
In Table 9, we also test our method in more advanced model Our results demonstrate that the performance of Antidote can be generalized to more advanced model.
Table 9: Evaluation of Antidote on Llama3-8B.
Methods Harmful Score Finetune Accuracy
SFT 80.30 42.40
Vaccine 77.50 36.90
RepNoise 78.30 41.40
Lisa 74.40 41.30
LDIFS 71.50 15.90
Antidote 71.20 39.00
5.4 Statistical and System Evaluation
Harmful embedding drift. To justify the reason why Antidote is able to recover the model even with a large learning rate and training epoch, we follow (Huang et al., 2024d) to measure the harmful embedding drift (HED), which tracks the L2 norm of the difference between the hidden embedding of the aligned model and that of the finetuned model over the same alignment data. We show in Figure 5 how different learning rates and training epochs affect this statistic. As shown, Antidote maintains the HED on a small scale, while the HED of other mitigation strategies escalates with the growth of learning rate and training epochs. Particularly, note that Antidote and SFT share the same two identical processes (and therefore the same HED) in their first two stages, but after the one-shot pruning in the post-fine-tuning stage, the HED of Antidote is significantly lower. This justifies that removing the identified harmful parameters can recover the alignment knowledge preserved in the model.
@@ 5.4 Statistical and System Evaluation
In summary, our finding justifies that Antidote can minimize the hidden embedding drift over the alignment data by pruning some specific harmful coordinates, and thereby mitigating the harmful embedding drift issue mentioned in (Huang et al., 2024d).
Output logit drift visualization. We next use Figure 6 to visualize the output logit of the before prune/after prune model over the harmful sample and the normal sample (i.e., GSM8K). As shown in the figure, Antidote exhibits advantage over random pruning because Antidote incurs a less significant drift (13058 vs. 22000) over GSM8K samples between the before-pruned and after-pruned model and similar drift (24469 vs 26172) over the harmful samples. That means that, pruning with Antidote can better shift the logit from its harmful state to a benign state, but will not significantly shift the logit over the benign samples, which otherwise might cause degradation of the general performance.
Figure 6: Visualization of output logit. Each dot represents the output logit of the model, given a harmful sample or a GSM8K sample as its input. For example, to generate a red point, we input a GSM8K sample into the before-prune model and extract its logit.
@@
5.5 Hyper-parameters Analysis and Ablation Study
@@
Table 12: Antidote with different ways to calculate Wanda score under different poison ratio $p$. Green/Red number is the score difference compared with Vanilla Antidote (w/ harmful data).
@@
Impact of the size of realignment dataset. Per Eq. (1), we calculate the harmful Wanda score with the average statistics over the realignment dataset (harmful dataset). It is interesting to see how different numbers of harmful samples included in the re-alignment dataset can affect defense performance. Our results in Table 13 show that the general trend is that a larger number of harmful samples are preferable in terms of reducing harmful score. This is understandable because a larger number of harmful samples can better reflect the real harmful distribution therefore resulting in a more precise identification of harmful parameters. Another finding is that 1k samples of harmful data seem to be performing well, and the benefit of further increasing the sample number diminishes. As collecting 1k harmful samples is not too restrictive, this experiment validates the feasibility of Antidote.
@@ 5.6 Extensions
5.7 Visualization
As follows, we show how different methods respond to a malicious prompt. We used the finetuned model over the default setting for evaluation. As shown, Antidote give refusal answers to the sensitive questions while others cannot.
@@ 6 Conclusion
Acknowledgment
This research is partially sponsored by the NSF CISE grants 2302720, 2312758, 2038029, an IBM faculty award, a grant from CISCO Edge AI program. This research is supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA. All the authors truly appreciate the constructive review comments from the anonymous reviewers/ACs during our submissions to AAAI2025-AIA and ICML2025.
Impact Statement
This paper studies a security vulnerability of the LLM finetune API, known as harmful fine-tuning attack. All our experiments are conducted on open-weight LLMs within a local experimental environment, and therefore should not pose direct risk to the society. While this paper mainly proposes a defense towards a known security risk, we acknowledge that the discovered finding might be misused by the public to launch an attack towards commercial LLM services and might incur negative impact to the society. Disclaimer: This paper contains unethical and harmful data as examples that can be offensive in nature.
@@ Appendix A Experiment Setup
A.1 Detailed Setup
Training hyper-parameters. For the alignment stage, we use the learning rate of 1e31e-3 and train for 20 epochs. For the fine-tuning stage, we use the default learning rate 1e41e-4 and train for 20 epochs in default. We follow (Huang et al., 2024d) to utilize the double LoRA implementation, i.e., for the alignment stage and fine-tuning stage, two different LoRA adaptors are used. The rank of LoRA adaptor is 256 with LoRA alpha set to 4. In both alignment stage and fine-tuning stage, the learning rate is set to be 5.
Prompt template. We follow (Huang et al., 2024d) to use the Alpaca prompt template (Taori et al., 2023) in the following for constructing supervised dataset for alignment/finetuning.